# IFT-UAM/CSIC-25-85

### v1

@@ VI Gravitational waves from superheating
We have identified the conditions under which a scalar $\phi$ can find itself in a metastable minimum at high-temperature. We have not committed to a particular realisation, but we believe our findings can be applied to many scenarios, for example to the Higgs during a modified electroweak phase transition Grojean:2004xa ; Delaunay:2007wb ; Ellis:2018mja ; Ellis:2019oqb ; Bruggisser:2018mus ; Espinosa:2007qk ; Beniwal:2017eik ; Barger:2007im ; Espinosa:2011ax ; Kozaczuk:2019pet ; Kurup:2017dzf ; Azatov:2022tii . We here highlight the main phenomenological consequence of gravitational wave production, leaving the technical details to future work.
The physics of the gravitational waves produced at the superheating phase transitions depends a lot on the details of the reheating sector of section IV.3. Here we compute only the contribution from the vacuum, in order to get estimates for the amplitude and the frequency at the peak. We follow Athron:2023xlk for the derivation.
Here we notice that the temperature raise happens when the universe finds itself in a phase of matter dominance - driven by the reheating sector - and just when $H(t)\approx\Gamma_{R}$ the universe is in radiation at global temperature $T_{R}$. [^4] This is the usual reheating temperature of the universe. Clearly $T_{\rm max}(t_{\rm max})\ll T_{R}$, and the system is also potentially subject to dilution if $t_{\rm max}$ is an epoch well before the start of radiation dominance.
@@
The inverse transition while heating happens certainly during matter dominance, since the temperature raise is present only for time scales much shorter than the decay rate $\Gamma_{R}^{-1}$. The inverse transition while cooling (after the system drops below $T_{\rm max}$ adiabatically), could be happen both during matter or radiation dominance, depending on whether the nucleation happens before or after $T_{R}$. The dilution factor $\eta$ for the inverse (while heating) is the fourth power of the ratio of the scale factor at nucleation and at $T_{R}$, leading to the estimate
@@
where we see that $\Omega_{\rm gw}\propto T_{n}^{8}/T_{R}^{8}$. Here $g_{*}$ is the relativistic number of degrees of freedom, $\Omega_{\gamma}$ and $T_{\gamma}$ the abundance and temperature of the photons today, while $v_{w}$ is the velocity of the bubble wall. The direct $\rightarrow$ transition is most likely to happen during radiation dominance and not subject (or not so much) to dilution, and it is given by the above amplitude with $\eta=1$.
The peak frequency of the gravitational wave spectrum is related to the value of $\beta/H$ and it is again sensitive to dilution for the inverse transitions. In general, we write it as
@@
In our model, the size of $\beta/H$ at the nucleation can be estimated as follows, recalling that $\beta\equiv(d\Gamma/dt)/\Gamma$. In our context the time evolution is not monotonic in terms of the evolution of the temperature of our sector. Therefore we expect new features in the value of $\beta/H]_{n}$, depends on the temperature evolution as
@@
The last term in parenthesis has different values for the different types of transitions. For direct transition while cooling, it equates to 1-1, while a different factor arises for the inverse transitions $(\leftarrow,h)$ and $(\leftarrow,c)$. In particular we find
@@
where $O(10)$ comes from the fact that the transition happens when $S_{3}/T$ approaches its asymptotic almost constant value. We see that the fact that the temperature of the sector is raising on time scales shorter than Hubble and $\Gamma_{R}$, makes $\beta/H$ effectively bigger by a factor $H_{*}/\Gamma_{R}$. To support our estimate we compute $d\log\Gamma/d\log T$ in figure 6. Here $d\log\Gamma/d\log T$ can be thought of as an intrinsic value for $\beta/H$ before the time dependence of the temperature is taken into account.
Figure 6: Comparison between the “intrinsic $\beta/H=d\log\Gamma/d\log T$” and the corresponding Euclidean bounce action. Solid curves denote tunnelling toward the symmetric vacuum ($\leftarrow$), while dashed curves denote tunnelling toward the broken vacuum ($\rightarrow$). The grey dashed horizontal line at $d\log\Gamma/d\log T=4$ indicates the minimal intrinsic value attained in the scale‐invariant regime.
@@ VII Conclusions
In this paper, we have introduced the concept of super‐heating in cosmological phase transitions. Our benchmark are sectors with a large number of bosons, with classical scale invariance, and we have identified the broad conditions that allows a metastable vacuum to persist up to arbitrarily high temperatures.
During super-heating (or just heating), these sectors have an effective potential that can give rise to an inverse first‐order phase transition upon heating ($\leftarrow$), followed inevitably by the corresponding direct transition during cooling ($\rightarrow$). The former towards a vacuum with massless states, the latter towards a vacuum with massive states.
All the phase transitions towards the origin ($\leftarrow$) are of inverse type, and their corresponding from the origin ($\rightarrow$) are direct. By computing the bounce action $S_{3}/T$, the nucleation criterion from Eq. (18) shows that even if an inverse phase transition while heating is not happening, it might happen during the subsequent cooling.
In our scenario the inverse transition induced by heating always takes place during a matter‑dominated era due to the slow heating, while the one upon cooling depends on the reheating model. If it occurs, then the corresponding direct transition while cooling inevitably follows. As noticed in Buen-Abad:2023hex , we find that even more generally, these pairs of phase transitions could generate a distinctive double‑peaked gravitational‑wave spectrum. The first peak, if arising in the matter dominated phase, receives additional redshift dilution, while the second peak originates from the familiar radiation‑dominated direct transition. Furthermore, by estimating the phase‑transition duration we demonstrate that a rapid temperature rise enhances $\beta/H$ compared to usual phase transitions. We have emphasised the importance of establishing thermal equilibrium well before the temperature where the tree-level instability disappears, and this is correlated with the prediction for gravitational waves.

### v2

@@ VI Gravitational waves from superheating
We have identified the conditions under which a scalar $\phi$ can find itself in a metastable minimum at high-temperature. We have not committed to a particular realisation, but we believe our findings can be applied to many scenarios, for example to the Higgs during a modified electroweak phase transition Grojean et al. (2005); Delaunay et al. (2008); Ellis et al. (2018, 2019); Bruggisser et al. (2018); Espinosa and Quiros (2007); Beniwal et al. (2017); Barger et al. (2008); Espinosa et al. (2012); Kozaczuk et al. (2020); Kurup and Perelstein (2017); Azatov et al. (2022). Here we highlight the main phenomenological consequence of gravitational wave production, leaving the technical details to future work.
The physics of the gravitational waves produced at the superheating phase transitions depends strongly on the details of the reheating sector of section IV.3. Here we compute only the contribution from the vacuum, in order to get estimates for the amplitude and the frequency at the peak. We follow Athron et al. (2024) for the derivation.
Here we notice that the temperature rise happens when the universe finds itself in a phase of matter dominance - driven by the reheating sector - and just when $H(t)\approx\Gamma_{R}$ the universe is in radiation at global temperature $T_{R}$. [^6] This is the usual reheating temperature of the universe. Clearly $T_{\rm max}(t_{\rm max})\ll T_{R}$, and the system is also potentially subject to dilution if $t_{\rm max}$ is an epoch well before the start of radiation dominance.
@@
The inverse transition while heating certainly happens during matter dominance, since the temperature raise is present only for time scales much shorter than the decay rate $\Gamma_{R}^{-1}$. The inverse transition while cooling (after the system drops below $T_{\rm max}$ adiabatically), could happen during either matter or radiation dominance, depending on whether the nucleation happens before or after $T_{R}$. The dilution factor $\eta$ for the inverse (while heating) is the fourth power of the ratio of the scale factor at nucleation and at $T_{R}$, leading to the estimate
@@
where we see that $\Omega_{\rm gw}\propto T_{n}^{8}/T_{R}^{8}$. Here $g_{*}$ is the relativistic number of degrees of freedom, $\Omega_{\gamma}$ and $T_{\gamma}$ are the abundance and temperature of the photons today, while $v_{w}$ is the velocity of the bubble wall. The direct $\rightarrow$ transition is most likely to happen during radiation dominance and not subject (or not so much) to dilution, and it is given by the above amplitude with $\eta=1$.
The peak frequency of the gravitational wave spectrum is related to the value of $\beta/H$, and it is again sensitive to dilution for the inverse transitions. In general, we write it as
@@
In our model, the size of $\beta/H$ at the nucleation can be estimated as follows, recalling that $\beta\equiv(d\Gamma/dt)/\Gamma$. In our context, the time evolution is not monotonic in terms of the evolution of the temperature of our sector. Therefore we expect new features in the value of $\beta/H]_{n}$, depends on the temperature evolution as
@@
The last term in parentheses has different values for the different types of transitions. For direct transition while cooling, it equates to 1-1, while a different factor arises for the inverse transitions $(\leftarrow,h)$ and $(\leftarrow,c)$. In particular, we find
@@
where $O(10)$ comes from the fact that the transition happens when $S_{3}/T$ approaches its asymptotic almost constant value. We see that the fact that the temperature of the sector is rising on time scales shorter than Hubble and $\Gamma_{R}$, makes $\beta/H$ effectively bigger by a factor $H_{*}/\Gamma_{R}$. To support our estimate we compute $d\log\Gamma/d\log T$ in figure 7. Here $d\log\Gamma/d\log T$ can be thought of as an intrinsic value for $\beta/H$ before the time dependence of the temperature is taken into account.
Figure 7: Comparison between the “intrinsic $\beta/H=d\log\Gamma/d\log T$” and the corresponding Euclidean bounce action. Solid curves denote nucleation toward the symmetric vacuum ($\leftarrow$), while dashed curves denote nucleation toward the broken vacuum ($\rightarrow$). The grey dashed horizontal line at $d\log\Gamma/d\log T=4$ indicates the minimal intrinsic value attained in the scale‐invariant regime.
@@ VII Conclusions
In this paper, we have introduced the concept of super‐heating in cosmological phase transitions. Our benchmarks are sectors with a large number of bosons, with classical scale invariance, and we have identified the broad conditions that allow a metastable vacuum to persist up to arbitrarily high temperatures.
During super-heating (or simply heating), these sectors have an effective potential that can give rise to an inverse first‐order phase transition upon heating ($\leftarrow$), followed inevitably by the corresponding direct transition during cooling ($\rightarrow$). The former proceeds towards a vacuum with massless states, the latter towards a vacuum with massive states.
All the phase transitions towards the origin ($\leftarrow$) are of inverse type, and their corresponding ones from the origin ($\rightarrow$) are direct. By computing the bounce action $S_{3}/T$, the nucleation criterion from Eq. (18) shows that even if an inverse phase transition while heating does not occur, it may happen during the subsequent cooling.
In our scenario, the inverse transition induced by heating always takes place during a matter‑dominated era due to the slow heating, while the one upon cooling depends on the reheating model. If it occurs, then the corresponding direct transition while cooling inevitably follows. As noticed in Buen-Abad et al. (2023), we find that even more generally, these pairs of phase transitions could generate a distinctive double‑peaked gravitational‑wave spectrum. The first peak, if arising in the matter dominated phase, receives additional redshift dilution, while the second peak originates from the familiar radiation‑dominated direct transition. Furthermore, by estimating the phase‑transition duration we demonstrate that a rapid temperature rise enhances $\beta/H$ compared to usual phase transitions. We have emphasised the importance of establishing thermal equilibrium well before the temperature where the tree-level instability disappears, and this is correlated with the prediction for gravitational waves.